OpenStudy (anonymous):

The sides of a square are increasing at a rate of 5 m/s. the area of the square is changing at a rate of (blank) m/s^2 when the sides are 30m long.

OpenStudy (anonymous):

Okay start with: \[ A=x^2 \]

OpenStudy (anonymous):

Can you implicitly differentiate this equation with respect to \(t\)?

OpenStudy (anonymous):

length = 5(t) + length of original side area = length^2 dl/dt = 5 da/dl = 2(L)

OpenStudy (anonymous):

da/dt = 10(L)

OpenStudy (anonymous):

Don't come up with any explicit equations of \(t\).

OpenStudy (anonymous):

dA/dx = 2x

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

I would think the sides are changing at a rate of 2(x) where x = 30 = length of side

OpenStudy (anonymous):

= 60

OpenStudy (anonymous):

\[ A'=2xx'=2(30)(5) \]

OpenStudy (anonymous):

oh because x is a function of time?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[ \frac{dA}{dt}=\frac{dA}{dx}\frac{dx}{dt} \]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

oh because x is a function of time?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Well you just have to remember that derivatives is all about which variable you differentiate.